For the FY 2018, Dorchester Company's balance sheet included the following current items: cash $29,000, accounts receivable $128,000, inventories $102,000, prepaid expenses $25,000, accounts payable $66,000, and accrued expenses $65,000.  Use this information to determine the Current Ratio. (Round & enter your answers to one decimal place.)
Answer #1

Current assets = Cash+Account receivable+Inventories+Prepaid expense = 29000+128000+102000+25000 = 284000

Current liabilities = 66000+65000 = 131000

Current ratio = Current assets/Current liabilities = 284000/131000 = 2.2